Angela is on her way from home from Jersey City into New York City for dinner. She walks 1 mile to the train station, takes the train 4/5 of the way, and takes a taxi 1/6 of the way to the restaurant. How far does Angela live from the restaurant?
Answer by Alan3354(69443)

Angela is on her way from home from Jersey City into New York City for dinner. She walks 1 mile to the train station, takes the train 4/5 of the way, and takes a taxi 1/6 of the way to the restaurant. How far does Angela live from the restaurant?
=========================
d = 1 + (4/5)d + (1/6)d
d = 1 + (29/30)d
d - 29d/30 = 1
d/30 = 1
d = 30 miles